OpenStudy (ffasinger):

What is the equation of the following line? Be sure to scroll down first to see all answer options.

OpenStudy (melody323):

What are the options?

Mehek (mehek14):

use the 2 points \(\dfrac{0-1}{0-(-3)}-?\)

OpenStudy (ffasinger):

A y=-x B y=-1/3 C y=x D y=3x E y=1/3x F y=-3x

OpenStudy (ffasinger):

@okdutchman7

Mehek (mehek14):

subtract what I put

OpenStudy (ffasinger):

@jabez177

OpenStudy (ffasinger):

i subtacted that but it did not come out right

Mehek (mehek14):

are you sure? what is 0 - 1? = -1 what is 0 - (-3)? = 3 -1/3 easy

OpenStudy (ffasinger):

so it woul be B

Mehek (mehek14):

correct